### v1

# SVLA: A Unified Speech-Vision-Language Assistant with Multimodal Reasoning and Speech Generation

Abstract
Large Vision-Language Models have shown impressive capabilities in tasks such as image captioning, visual question answering, and cross-modal retrieval. However, there are still significant challenges that need to be addressed in order to fully unlock the potential of these models. First, integrating speech, text, and vision into a unified model is particularly difficult for tasks like Spoken Image Captioning and Spoken Visual Question Answering, where the interaction between these modalities introduces additional complexity. Second, existing speech generation approaches differ—some generate speech directly, while others use an intermediate text step—but their impact on fluency, coherence, and accuracy remains unexplored. Third, current speech evaluation relies heavily on Automatic Speech Recognition, which introduces biases and fails to directly assess speech quality. To address these challenges, we propose SVLA, a unified Speech-Vision-Language Assistant based on a decoder-only transformer architecture that seamlessly integrates multimodal inputs and outputs. We enhance model performance with a large-scale speech-text-image dataset containing 38.2 million examples and 64.1 hours of TTS-generated speech. Additionally, we introduce Speech-VQA Accuracy, a novel acoustic evaluation metric for assessing speech-based VQA tasks. Our approach advances multimodal understanding and generation, facilitating more effective integration of speech, text, and vision (http://github.com/vlm-svla/svla).
1 Introduction
Large Language Models (LLMs) like LLaMA [1], GPT-3 [2], InternLM-2 [3], and Qwen [4] excel in textual processing but are limited by their unimodal design. To address real-world needs, Large Vision-Language Models (LVLMs) such as GPT-4V [5], Gemini [6], BLIP-2 [7], Llava [8], and Flamingo [9] integrate visual data, while Unified Speech-Text-Language Models further incorporate speech to enhance accessibility and natural interactions. Recent efforts—exemplified by closed-source ChatGPT-4o [10] and open-source models like LLaMA 3.1 [11], AnyGPT [12], and NextGPT [13]—have expanded multimodal capabilities to speech-vision tasks, such as generating images from speech or creating music from visual inputs.
However, despite the promising advancements made so far, significant challenges remain that must be addressed in order to fully unlock the potential of these models. First, these models lack robust reasoning abilities for tasks like image captioning (IC) and visual question answering (VQA), where inputs may be text or speech and outputs may be speech or text, limiting their applicability to complex multimodal interactions. Second, existing speech generation approaches vary, with some models generating speech directly, while others use text instructions as an intermediary step. However, no prior research has systematically compared these two techniques. This distinction is crucial, as the choice between these paradigms significantly impacts fluency, coherence, and response accuracy in multimodal AI. For example, when asked “What are the people doing?”, a text-instruction model first generates a structured textual response before generating speech tokens (e.g., “The textual answer is ‘running.’ Therefore, the speech answer is ‘running.’(speaking)”). While this ensures consistency between text and speech outputs, it introduces redundancy and unnatural phrasing, which may not align with natural spoken communication. In contrast, direct speech generation bypasses the textual constraint, producing more fluid spoken responses but at the risk of greater variability, pronunciation inconsistencies, or a loss of linguistic structure. Understanding these trade-offs is essential for advancing multimodal reasoning, spoken language generation, and user interaction quality in AI systems. Finally, existing evaluation metrics for multimodal learning focus mainly on text and vision, offering limited insight into speech-based tasks.
To address the challenges mentioned above, we introduce SVLA, a Speech-Vision-Language Assistant—a unified multimodal system designed for end-to-end reasoning and generation. A speech discretization tokenizer converts raw speech into discrete semantic tokens, enabling the model to process speech similarly to text. This design allows a traditional generative LLM to seamlessly handle tri-modal inputs—speech, text, and images—and generate outputs in either speech or text. SVLA supports a broad range of tasks, including ASR, TTS, Image Captioning (IC), and Visual Question Answering (VQA). Importantly, the system is flexible in handling various modality combinations. For example, it can generate text outputs from speech-based inputs (speech-to-text tasks) or produce speech outputs from text-based inputs (text-to-speech tasks). To evaluate SVLA, we propose Acoustic Speech-VQA Accuracy, a novel metric that directly assesses spoken responses—bypassing ASR conversion—providing a more accurate and unbiased evaluation of the model’s spoken language capabilities. Additionally, SVLA is assessed on a new, large-scale speech-text-vision dataset that empowers models to perform visual tasks using both text and speech. This dataset is built by integrating several widely-used multimodal and speech resources, including VQAv2[14], LAION-600M[15], Visual Genome[16], LibriHeavy[17], LibriSpeech[18], CommonVoice[19], A-OKVQA[20], VizWiz[21], GQA[22], and COCO-Caption[23]. To ensure natural and diverse speech data, we utilize a publicly available TTS system[24] that adjusts accent and speed, enriching linguistic variety and prosody.
Our contributions are as follows:
  • We introduce SVLA, a unified multimodal assistant designed for end-to-end speech and text reasoning in visual understanding.
  • We present a large-scale speech-text-vision dataset built from widely used multimodal and speech benchmarks, providing aligned text and speech annotations to enhance multimodal learning and facilitate speech-vision-language tasks in both textual and spoken formats.
  • We propose Acoustic Speech-VQA Accuracy, a novel metric for evaluating spoken VQA responses without relying on ASR conversion, ensuring a more direct and unbiased assessment of speech-based reasoning.
  • We establish a comprehensive evaluation framework that systematically benchmarks models across diverse input-output configurations.
2 Related Works
Below, we explore key advancements in integrating speech into multimodal LLMs, compare direct and instruction-based speech generation approaches, and examine evaluation metrics for speech in multimodal tasks.
Speech Multi-modal Large Language Models: Recent efforts have focused on integrating audio capabilities into LVLMs to enhance multimodal interactions. Early models supported only speech input, limiting their effectiveness in conversational AI [25, 26]. To address this, some approaches incorporated TTS for spoken responses [27, 11], but these pipelines introduce higher computational costs, inference latency, and limited contextual adaptation. OpenAI’s GPT-4o [28] overcomes these limitations by generating speech, enabling real-time, expressive interactions with preserved intonation and emotional cues. However, GPT-4o remains proprietary, prompting open-source efforts to replicate its multimodal capabilities. Models such as Next-GPT [13], Codi-1 [29], Codi-2 [30], UIO [31], and AnyGPT [12] have attempted vision-speech integration using pre-trained encoders. They mostly pay attend on spoken tasks as: Image Generation, Music Generation, Audio Captioning, Video Generation, ASR, and TTS . Yet, most remain limited to IC and VQA within a TTT paradigm, lacking support for crucial speech-vision tasks such as TTS, STT, and STS. However, TMT [32] acttually build a model to speak image captions, but is designed to translate between modalities one at a time, meaning it can handle one input modality at a time and then generate an output in another modality.
Speech-Text-Vision Datasets: While multimodal datasets like VQAv2 [14], LAION-600M [15], and Visual Genome [16] focus on text-vision alignment, they lack natural speech data, limiting their applicability to speech-vision reasoning. Some datasets incorporate audio-visual elements [33, 34, 35, 36], but audio is often restricted to environmental sounds or extracted speech, with questions presented in text. Speech-COCO [37] extends image captioning with speech annotations but remains small in scale and lacks diverse multimodal reasoning tasks. Moreover, existing datasets do not support seamless switching between speech and text, preventing models from dynamically adapting to user intent, such as generating spoken answers or textual captions based on instructions.
Direct vs. Instruction-Based Speech Generation:
Speech generation in multimodal AI follows two main approaches: direct speech generation and instruction-based speech generation. Direct models generate spoken responses end-to-end without intermediate text, while instruction-based models—such as ChatGPT-4o [10], AnyGPT, and Next-GPT [13, 29]—produce both text and speech simultaneously, with text guiding the spoken response. These models output structured text (e.g., ”The answer is ‘running’.”) before generating speech, ensuring consistency but introducing redundancy and unnatural phrasing. While instruction-based generation aligns text and speech outputs, it is less suited for natural conversations. However, a systematic comparison between these approaches remains unexplored, leaving open questions about their effectiveness for context-aware spoken interactions in multimodal AI.
Speech-Based Evaluation Metrics for Multimodal Tasks:
Evaluating text output is straightforward using established metrics like BLEU [38], ROUGE [39], Accuracy, and CIDEr [40]. In contrast, speech evaluation is challenging due to its continuous nature, prosody, and pronunciation variations, which traditional text-based metrics cannot fully capture. A common approach transcribes speech using ASR before applying text metrics [12, 13], but ASR errors can introduce inconsistencies and bias evaluations. Pretrained speech models like WavLM-TDNN [41] have been used to measure similarity in speech output tasks, but they do not account for reasoning quality or semantic correctness. This highlights the need for a dedicated evaluation metric to assessing spoken responses in multimodal reasoning tasks.
3 Data Generation
Figure 1: Pre-training data distribution across three modalities: Image, Text, and Speech.
A variety of image-text datasets—such as COCO-caption [23], and LAION-COCO [15]—offer abundant and diverse visual and linguistic data. We leverage these resources to build a robust, multi-purpose dataset by pairing each image with its corresponding text and converting that text into speech using the Melon-TTS model [24]. To ensure the generated audio sounds natural, we vary both the speaking speed (ranging from 0.7 to 1.3) and the accents (British, American, Indian, and Australian). Additionally, all audio is standardized at a 16 kHz sampling rate, striking a good balance between quality and size, and maintaining consistent clarity across all recordings.
3.1 Pre-train dataset
Our pre-training strategy is designed to develop a unified speech-vision-text model capable of handling TTS, ASR, Image Captioning (IC), and VQA. The training data is divided into two key subsets: (1) text-speech data, sourced from LibriHeavy [17], comprising 8M audiobook-transcript pairs to enhance TTS expressiveness and ASR accuracy; and (2) vision-text-speech data, which integrates LAION-COCO [15] (8M image-text pairs) and Visual Genome (VG) [16] to improve the model’s ability to interpret and respond to visual queries. To ensure seamless modality switching, we incorporate instructional prompts such as “Answer this question in speech” or “Write down what you see in the image”. Figure 1 shows the token distribution across three modalities: Image, Text, and Speech [^1] The number of text tokens is computed using Qwen-2.5-1.5B [42], while speech tokens are counted using SpeechTokenizer [43]. Image tokens are estimated at 256 tokens per image.. Additional details on pre-training are provided in Appendix A.
Figure 2: SVLA Architecture
3.2 Visual Instruction dataset
For supervised fine-tuning (SFT), we follow a similar structure but introduce multi-turn conversations to enhance coherence and context retention. The text-speech subset includes LibriSpeech [18] and CommandVoice [19] for diverse linguistic and command-based tasks. The vision-text-speech subset incorporates VQAv2, A-OKVQA [20], GQA [22], VizWiz [21], and COCO-Caption-2014, ensuring broad multimodal coverage. Unlike pre-training, SFT dialogues extend up to 4096 tokens, allowing the model to sustain long-form interactions. Modality-switching prompts guide the model’s transitions between speech, text, and vision tasks, ensuring flexible and adaptive multimodal reasoning. Further details on the SFT dataset can be found in Appendix A.
4 Model Architecture
Our architecture builds on LLaVA’s design [8] with additional components for speech processing. Our model integrates speech and vision inputs into an LLM, denoted as $f_{\theta}(x)$, where $\theta$ represents the model parameters. The overall network architecture is depicted in Figure 2.
4.1 Vision Encoder
Given an input image $X_{v}$, we employ a pretrained VIT-based vision encoder, such as CLIP[44], denoted as $g(\cdot)$, to extract high-level visual representations:
$Z_{v}=g(X_{v}),\quad Z_{v}\in\mathbb{R}^{n_{v}\times d_{v}}.$ (1)
Here, $Z_{v}$ is obtained from the final layer of the vision encoder, where $n_{v}$ is the number of image patches. Each visual token in $Z_{v}$ has a feature dimension of $d_{v}$.
Next, to align visual embeddings with the LLM’s word embedding space, we apply a learnable linear projection layer $W_{v}$:
$H_{v}=W_{v}Z_{v},\quad H_{v}\in\mathbb{R}^{n_{v}\times d_{h}}.$ (2)
Here, $H_{v}$ represents the projected visual tokens, now residing in the same dimensional space $d_{h}$ as the LLM’s text embeddings, ensuring compatibility for multimodal fusion.
4.2 Speech Encoder and Tokenization
For speech input $X_{s}$, a pretrained speech encoder $s(\cdot)$ generates a sequence of discrete tokens:
$Z_{s}=\{q_{1},q_{2},\dots,q_{T_{s}}\},\quad q_{i}\in\{1,2,\dots,d_{s}\},$
[…]
4.3 Multimodal Fusion
Given textual tokens $Z_{t}$ from input text $X_{t}$, discrete speech tokens $Z_{s}$, and visual embeddings $H_{v}$, we construct the multimodal token sequence:
$X=[bos,\,boi,\,img,\,eoi,\,Z_{t},\,boa,\,Z_{s},\,eoa,\,eos],\quad X\in\mathbb{R}^{n}$
where $n$ is the initial token sequence length. Special tokens ($bos,boi,eoi,boa,eoa,eos$) explicitly mark the beginning and end of the sequence, image tokens, and audio tokens.
After embedding, we replace the placeholder embedding of the token $img$ with the visual embeddings $H_{v}$:
$E^{\prime}=[E_{bos},\,E_{boi},\,H_{v},\,E_{eoi},\,E_{Z_{t}},\,E_{boa},\,E_{Z_{s}},\,E_{eoa},\,E_{eos}],$
$\quad E^{\prime}\in\mathbb{R}^{n^{\prime}\times d_{h}}$, where $n^{\prime}=n_{v}+n-1$. Finally, the backbone LLM processes $E^{\prime}$ to achieve multimodal understanding.
4.4 Speech Decoding
The multimodal LLM outputs a sequence of predicted tokens, from which the speech-related tokens $Z_{s}^{\prime}$ are extracted between the special tokens $boa$ and $eoa$. A pretrained speech decoder $s^{-1}(\cdot)$ then transforms these discrete tokens back into speech waveforms:
$X_{s}^{\prime}=s^{-1}(Z_{s}^{\prime}),$
where $Z_{s}^{\prime}$ represents the predicted discrete speech token sequence, and $X_{s}^{\prime}$ is the resulting synthesized speech audio.
Table 1: Comparison of models on Image Captioning and VQA tasks. We evaluate Image Captioning on COCO-Caption-2014, COCO-Caption-2017, and Flickr8k datasets, while VQA performance is assessed on VQAv2-val, OKVQA-test, GQA-test, and VizWiz datasets. Modalities are denoted as I (Image), T (Text), and S (Speech). * indicates results on the test-dev set.
Model Backbone Input $\rightarrow$ Ouput Image Captioning VQA
COCO-2014-test COCO-2017-test Flickr8k VQAv2-val OKVQA-test GQA-test VizWiz
[…]
Table 2: Speech-Based Captioning and VQA Performance Comparison
Model Input $\rightarrow$ Ouput Image Captioning (Similarity) VQA (Speech-VQA-Accuracy)
COCO-2014 COCO-2017 Flickr30k VQAv2-val OKVQA-test GQA-test VizWiz
[…]
4.5 Text-Instruction Outputs
During the pre-train stages (can be found in Appendix A.2), the model generate text tokens and speech tokens directly. Similarly, during the SFT stage, the model generates responses directly in textual form. However, for speech output, we use a structured prompting system to guide the model in reasoning and generating coherent spoken responses. Unlike text generation, speech output introduces additional complexities, requiring structured guidance to ensure fluency, coherence, and semantic alignment. For ASR and TTS tasks, we use simple yet effective prompts to map between speech and text. ASR follows “The transcript of the given speech is:” to extract text, while TTS uses “This is how your text sounds in speech:” to generate speech. For speech-vision tasks, where the model must generate speech from visual inputs, we adopt an instruction-based speech generation approach. Here, textual captions and textual answers serve as intermediate reasoning steps to align the model’s understanding before producing spoken outputs. In image captioning, we explicitly structure the prompt as “The textual caption is ‘caption’. Therefore, the audio caption is:” followed by discrete speech tokens. This ensures that the speech output remains faithful to the structured textual representation. Similarly, for VQA tasks, we prompt the model with “The textual answer is ‘answer’. Therefore, the audio answer is:” before generating discrete speech tokens. This method enforces a structured reasoning process, ensuring that the model first forms a grounded textual response before converting it into speech.
5 Experiments
This section details our experimental setup, the metrics used to assess performance, and the results achieved.
5.1 Implementation Details
We use Qwen2.5-1.5B as our backbone LLM and train it with PyTorch. The vision encoder is CLIP-large-patch14-336 [44], which producing $256$ total tokens for each image. Speech data is handled by the SpeechTokenizer[43], which encodes each 1-second segment of audio into 50 discrete tokens; when advanced speech fidelity is required, SoundStorm-SpeechTokenizer [^2] https://github.com/ZhangXInFD/soundstorm-speechtokenizerextends the same quantization principle to more nuanced tasks. The weights of these speech components are from AnyGPT[12]. More complementary details can be found in Appendix B.
5.2 Metrics
Text-Output tasks:
For tasks where the model generates text outputs, we use Word Error Rate (WER) for ASR, CIDEr [40] for image captioning, and accuracy for VQA. These metrics provide a standardized evaluation framework for assessing performance across text-output tasks.
Speech-output tasks:
We evaluate speech generation using two methods: We transcribe model-generated speech with Whisper Medium [48] and compare the resulting text to human references (as in text-output tasks). However, ASR models can exhibit biases that introduce transcription errors, thereby distorting the perceived quality of the generated speech. Therefore, we use WavLM-TDNN [41] to extract speech embeddings from both generated and reference speech and measure their similarity (e.g., via cosine similarity). This directly compares acoustic properties without relying on ASR. Apart from that, we propose a Speech-Vqa-Accuracy that incorporates the above similarity. Given a generated answer $\hat{A}$ and $N$ human-annotated references $A_{1},A_{2},\dots,A_{N}$, let the set of unique references be $\{U_{1},U_{2},\dots,U_{n}\}$ (where $n\leq N$). We compute the similarity $\mathrm{sim}(\hat{A},U_{j})$ for each $U_{j}$, and choose the closest reference:
$U_{\mathrm{best}}=\arg\max_{U_{j}}\,\mathrm{sim}(\hat{A},U_{j}).$ (3)
Let $\mathrm{count}(U_{\mathrm{best}})$ be the number of times $U_{\mathrm{best}}$ appears in the original $N$ references. We define a weight:
$W(U_{\mathrm{best}})=\begin{cases}1,&\text{if }\mathrm{count}(U_{\mathrm{best}})\geq 0.3N,\\ \tfrac{2}{3},&\text{if }\mathrm{count}(U_{\mathrm{best}})=0.2N,\\ \tfrac{1}{3},&\text{if }\mathrm{count}(U_{\mathrm{best}})=0.1N.\end{cases}$ (4)
Finally, we define the Speech-VQA Score:
$\mathrm{Score}_{\mathrm{VQA}}=I\!\bigl{(}\mathrm{sim}(\hat{A},U_{\mathrm{best}})\,\geq\,\tau\bigr{)}\,\cdot\,W(U_{\mathrm{best}}),$ (5)
where $\tau$ is a similarity threshold, and $I(\cdot)$ is an indicator function that is 1 if the condition is true, and 0 otherwise. In this paper, we set $\tau$ to 0.7 based on empirical observations that this threshold best balances precision and recall, effectively distinguishing perceptually similar speech segments from dissimilar ones.
5.3 Results
Figure 3: Visualization of the model’s attention maps when aligning image features with text (and/or spoken) queries.
Table 3: Performance of Text-Speech Tasks. We evaluate ASR on LibriSpeech [18] test-clean and TTS on VCTK [49]
Model Backbone ASR TTS
LibriSpeech VCTK
WER WER Similarity
Human-level - 5.8 1.9 0.93
Wav2vec 2.0 [38] - 2.7 - -
Whisper Large V2 [48] - 2.7 - -
VALL-E [50] - - 7.9 0.75
VILA [51] Mixtral 8x7B [52] 8.1 - -
AnyGPT-7B LLaMA-2-7B 8.5 8.5 0.77
MIO-Ins [47] Yi-6B-Base [53] 10.3 4.2 -
SVLA-2B Qwen-1.5B 10.2 21.7 0.69
SVLA-2B-Text-Ins Qwen-1.5B 9.2 11.2 0.74
Text-Speech Performance:
From Table 3, SVLA-2B-Text-Ins outperforms its direct-speech counterpart, SVLA-2B, on both ASR and TTS tasks. By prompting the model with structured text (as “This is the transcript:” and “This is how the text sounds:”), it reduces ASR WER from 10.2 to 9.2 and TTS WER from 21.7 to 11.2, and increases similarity from 0.63 to 0.74. Although both models still lag behind specialized ASR systems like Whisper Large V2 and Wav2vec 2.0 (each with WER 2.7), SVLA-2B-Text-Ins is competitive with the larger AnyGPT-8B (ASR WER 8.5, TTS WER 8.5, similarity 0.77). Direct speech generation (SVLA-2B) avoids verbose prompts but can sacrifice clarity, whereas instruction-based speech generation (SVLA-2B-Text-Ins) leverages explicit textual guides to improve transcription and synthesis accuracy, demonstrating the trade-off between naturalness and performance.
Image Captioning Performance:
From Table 1, instruction tuning in SVLA-2B-Text-Ins has minimal impact on text-only performance, evidenced by near-identical CIDEr scores on COCO-2014 (120.0 vs. 120.2) relative to the non-instruction version, SVLA-2B. However, when generating spoken captions, SVLA-2B scores close to zero, whereas SVLA-2B-Text-Ins jumps to around 56—an improvement driven by structured prompts (as “The textual caption is … Therefore, the audio caption is: …”) that significantly boost speech accuracy. This advantage is also clear in acoustic similarity metrics: on COCO-2014, SVLA-2B achieves 79.7, while SVLA-2B-Text-Ins rises to 89.3, reflecting more faithful speech outputs (Table 2). While TMT attains an even higher speech-only result (78.7 of CIDEr scores on COCO) by treating each modality as a separate “language”, it handles only one input modality at once and thus lacks a unified approach to image+text or image+speech inputs. Moreover, Next-GPT-7B surpasses SVLA-2B in text tasks (158.3 vs.  120 CIDEr) but does so at a larger 7B scale, while SVLA-2B (Qwen2.5-1.5B) maintains parameter efficiency and handles both text and speech.
VQA Performance:
In Table 1 shows that SVLA-2B-Text-Ins stands out in the I+T → T VQA task by delivering competitive performance despite utilizing a much smaller Qwen-1.5B backbone compared to the 7B-sized alternatives. While models like LLaVA-1.5 and Next-GPT-7B achieve strong scores on benchmarks such as VQAv2, OKVQA, and GQA, our SVLA-2B-Text-Ins variant demonstrates impressive results with a VQAv2 score of 69.7 and a VizWiz accuracy of 58.0, rivaling or even surpassing larger models. On the other hand, the model shows a significant drop—by approximately 20% in OKVQA and GQA—when evaluated on the I+S → T VQA task, indicating that the model performs better when the question is provided in text form. Notably, similar to the image captioning task, the instructive output strategy does not substantially impact performance in text-output VQA tasks, as the accuracy difference between non-instruct and instruct configurations is minimal.
The performance of speech outputs is generally worse than that of text outputs by roughly 15–20% on VQA tasks. In the ASR-based evaluation (Table1), the SVLA-2B-Text-Ins model operating in the I+T → S configuration achieves a VQAv2 score of 37% and an OKVQA score of 11.6% , compared to 29.4% and 6.08% , respectively, in the I+S → S configuration. This discrepancy indicates that integrating a text input—as opposed to directly processing speech—preserves richer semantic details. The acoustic Speech-VQA accuracy evaluation further highlights this benefit, with the SVLA-2B-Text-Ins model recording a VQAv2 accuracy of 51.7% and an OKVQA accuracy of 69.6% in I+T → S, versus 44.5% and 62.5% in I+S → S. These improvements are likely due to the structured intermediate text-generation stage, which reduces the ambiguity inherent in raw speech processing and results in more semantically faithful responses for complex VQA tasks.
5.4 Ablation
Where does the model look to the image when writing and speaking:
To investigate how the model allocates visual attention during text and speech generation, we visualize the attention patches corresponding to specific words in both modalities in Figure 3. In these examples, the text yields more focused attention than the speech. In Case A, text attention precisely highlights both individuals for “people” and localizes the “frisbee”, while speech attention is more diffused. In Case B, text cleanly isolates the rider for “man” and the wheels for “skateboard”, whereas speech spreads across a wider area. In Case C, text remains sharply concentrated on the rider for “man” and the animal for “horse”, but speech again shows broader contextual focus. This precision in text helps achieve better performance in visual tasks, as it zeroes in on the exact objects needed for accurate descriptions.
Effect of Similarity Threshold on Speech-based VQA
Figure 4: Impact of Similarity Threshold on acoustic Speech-VQA Accuracy.
(a) SVLA-2B-Text-Ins
(b) SVLA-2B
Figure 5: Comparison of SVLA-2B-Text-Ins and SVLA-2B in Multimodal Image Captioning and VQA Responses.
In Figure 4, we examine how different similarity thresholds (0.6 to 0.9) affect speech-based VQA accuracy. As the threshold increases, all models exhibit a sharp decline in accuracy, underscoring the sensitivity of speech-based metrics to acoustic matching. Notably, instruction-tuned models consistently outperform their non-instruction-tuned counterparts across all thresholds, suggesting that structured guidance improves both the semantic and acoustic quality of the generated speech. This result also highlights the importance of threshold selection: a higher threshold imposes stricter acoustic requirements, potentially penalizing answers that are semantically correct yet acoustically divergent from the references.
Does Text Guidance Make Speech Smarter?
According to Figure 5, SVLA-2B-Text-Ins generates more structured and coherent speech than SVLA-2B, enhancing semantic alignment and reasoning. By first generating a structured textual response, the model ensures that speech output follows a well-defined linguistic structure. This intermediate step helps maintain consistency, reduces ambiguity, and improves multimodal alignment, leading to clearer and more contextually relevant spoken responses. In contrast, SVLA-2B, which generates speech directly, tends to produce more random and less structured outputs due to the absence of textual guidance. This approach increases the likelihood of semantic misinterpretation, prosody misalignment, and overall incoherence.
The Limits of ASR-Based Evaluation:
ASR-based evaluation falls short in assessing speech quality because it misinterprets minor phonetic variations as errors, even when the spoken response is nearly correct. As seen in Figure 5(a), the model’s speech output closely follows the intended text, yet ASR misrecognizes words like “camera” as “camaral” and “kimeroh” or “tan” as “tanh”. These small differences, though perceptually acceptable, lead to unfair penalization since ASR focuses on exact text matching rather than acoustic similarity. This limitation reveals the need for an evaluation method that directly measures speech waveform similarity rather than relying on error-prone transcription.
6 Conclusion
In this work, we introduce SVLA, a unified Speech-Text-Vision Assistant capable of handling both language tasks (ASR, TTS) and vision-language reasoning tasks (image captioning, VQA). To support the community in building similar models, we also release a large-scale tri-modal dataset encompassing speech, text, and vision. Additionally, we analyze two settings for speech generation: directly producing spoken output and using a text prompt to guide speech synthesis. We further propose Speech-VQA-Accuracy, a new metric based on acoustic similarity for evaluating spoken VQA responses. Our experiments show that while the model performs better with text outputs, speech outputs benefit from an instructive text prompt, yielding more coherent.
Our study has several limitations. First, the speech data in our dataset is generated using a TTS model, which may lack the natural emotional and prosodic variations found in real human speech. Second, our architecture relies on a straightforward concatenation-based fusion, whereas more sophisticated approaches—such as Q-Former [7] or Perceiver Resampler [9]—could potentially yield better performance. Finally, due to resource constraints, we build on Qwen2.5-1.5B, which limits the model’s capacity. We hypothesize that employing a larger backbone would further enhance speech-related capabilities.
Appendix A Training
A.1 Data Generation
Dataset Task No. Sample Image Text Tokens Speech Tokens Speech Duration (s) Speech Duration (h)
Pretrain Dataset - Stage 1
Text-Speech tasks
Libriheavy TTS 1.0M 0 56.7M 730M 14.6M 4.1K
ASR 1.0M 0 60.0M 736M 14.7M 4.1K
Total - 2.0M 0 1.5M 7.4B 29.3M 8.2K
Pretrain Dataset - Stage 2
Text-Speech tasks
Libriheavy TTS 3.0M 0 151.6M 2.2B 44.2M 12.3K
ASR 3.0M 0 207.6M 2.2B 44.4M 12.3K
Image-Text-Speech tasks
Laion IC-TTT 5.8M 5.8M 111.1M 0 0 0
IC-TTS 5.8M 48.4M 1.1B 22.0M 6.1K
IC-STT 5.8M 63.1M 1.0B 20.2M 5.6K
IC-STS 5.8M 0 1.9B 38.6M 10.7K
VG VQA-TTT 1.3M 108K 12.1M 0 0 0
VQA-TTS 1.3M 19.1M 97.0M 1.9M 538
VQA-STT 1.3M 2.9M 321.0M 6.4M 1.8K
VQA-STS 1.3M 0 249.6M 182.4M 1.4K
Total - 34.3M 23.5M 615.9M 9.1B 182.8M 50.8K
Table 1: Statistics of the pretraining dataset
Dataset Task No. Sample Image Text Tokens Speech Tokens Speech Duration (s) Speech Duration (h)
Text-Speech tasks
Librispeech ASR 150K 0 5.4M 89.2M 1.8M 496.0
CommandVoice TTS 388K 0 4.8M 101.8M 2.0M 559.3
Image-Text-Speech tasks
VQA VQA-TTT 84K 83K 4.7M 0 0 0
VQA-TTS 42K 3.0M 13M 260K 72
VQA-STT 42K 282K 46.1M 921K 256
VQA-STS 84K 0 86.0M 17.2M 476
A-OKVQA VQA-TTT 50K 50K 520.0K 0 0 0
VQA-TTS 25K  [^3] K 1.4M 28K 8
VQA-STT 25K 31K 5M 102K 28
VQA-STS 50K 0 9.5M 190K 53
GQA VQA-TTT 72K 72K 11.5M 0 0 0
VQA-TTS 36K 6.7M 26.8M 536K 149
VQA-STT 36K 528.5K 100.5M 2.0M 558
VQA-STS 72K 0 168.0M 3.4M 933
VizWiz VQA-TTT 20K 20K 780.5K 0 0 0
VQA-TTS 10K 413K 883K 17.7K 5
VQA-STT 10K 28.8K 5.7M 114.0K 31.7
VQA-STS 20K 0 11.5M 230.7K 64.1
COCO-Caption-2014 IC-TTT 414K 83K 107.0M 0 0 0
IC-TTS 212K 3.3M 39.4M 789K 219
IC-STT 212K 2.4M 51.3M 1M 285
IC-STS 414K 0 163.7M 3.3M 909
Total - 2.5M 308K 152M 920M 33.9M 5102
Table 2: Statistics of the SFT dataset
Table 1 provides a detailed breakdown of the pretraining dataset, categorized into two stages: Stage 1, which consists solely of text-speech tasks, and Stage 2, which expands to include image-text-speech tasks. In Stage 1, the dataset comprises 2.0M samples from the Libriheavy corpus for TTS and ASR tasks, yielding a total of 1.5M text tokens, 7.4B speech tokens, and approximately 8.2K hours of speech. Stage 2 significantly scales up the dataset, incorporating Libriheavy (6M samples) alongside multimodal datasets such as Laion and VG, covering image-captioning (IC) and VQA tasks in both text and speech modalities. The total dataset spans 34.3M samples, 615.9M text tokens, 9.1B speech tokens, and 50.8K hours of speech, making it one of the most extensive speech-text-vision corpora for multimodal learning. Notably, the dataset supports a diverse set of multimodal tasks, including IC-TTT, IC-TTS, IC-STT, IC-STS, VQA-TTT, VQA-TTS, VQA-STT, and VQA-STS, ensuring broad coverage across different input-output combinations.
Table 2 presents the SFT dataset, covering both text-speech and image-text-speech tasks. The dataset includes 2.5M samples, with 308K image samples, 152M text tokens, and 920M speech tokens, totaling 33.9M seconds (5102 hours) of speech data. The text-speech tasks include 150K ASR samples from Librispeech and 388K TTS samples from CommandVoice, contributing 496 and 559 hours of speech, respectively. In the image-text-speech tasks, various VQA datasets (VQA, A-OKVQA, GQA, VizWiz) and COCO-Caption-2014 support text-based (VQA-TTT, IC-TTT), text-to-speech (VQA-TTS, IC-TTS), speech-to-text (VQA-STT, IC-STT), and speech-to-speech (VQA-STS, IC-STS) tasks. The dataset is diverse and well-balanced, ensuring broad multimodal coverage for fine-tuning models on speech, text, and vision-related tasks. Figure 1, Figure 2, Figure 3, and Figure 4 show the prompts of tasks used to train the models.
ASR (Automatic Speech Recognition) Prompts: “Please convert this audio to text: ” “Transcribe the following audio file, please: ” “Can you convert this speech to text? ” “Generate text from this audio recording: ” “Please write out what’s being said in this audio: ” “Turn this voice recording into text, please: ” “Please create a transcript of this audio: ” “Can you transcribe this audio? ” “Convert this spoken content into written text: ” “Please extract text from this speech: ” “Transcribe the spoken words in this audio file: ” “Create a written version of this audio: ” “Convert the spoken words to text: ” “Please generate a transcript from this recording: ” “Transform the audio into a text document: ” Example Usage:
"Please convert this audio to text: <speech_start>{speech_tokens}<speech_end>"
Figure 1: Prompts for ASR Tasks.
TTS (Text-to-Speech) Prompts: “Please convert this text to speech: ” “Turn this text into audio, please: ” “Generate speech from this text: ” “Please speak out this text: ” “Convert these words to speech, please: ” “Please make an audio version of this text: ” “Can you read this text aloud: ” “Transform this text into speech: ” “Please give a voice to this text: ” “Read out this text, please: ” “Create a spoken version of this text: ” “Convert the written text to speech: ” “Turn the following words into sound: ” “Provide an audio rendition of this text: ” “Generate an audio file of these words: ” Example Usage:
"Turn this text into audio: {transcript}"
Figure 2: Prompts for TTS Tasks.
Captioning Prompts: IC_TTT (Text-to-Text) and Caption_STS (Speech-to-Speech): “What do you see in the image?” “Explain what is shown in the picture.” “Provide a caption for the image.” “Describe the objects or people in the image.” “Tell what is happening in the picture.” “Explain the scene in the image.” “Describe the setting of the image.” “List the elements in the picture.” “What is the main focus of the image?” Example Usage:
[…]
  IC_TTS (Text-to-Speech): “Can you describe the image out loud?” “Read the description of the image aloud.” “Turn the image caption into spoken words.” “Provide a spoken description of the picture.” “Talk about what is shown in the image.” “Can you vocalize the image details?” “Use speech to explain the image.” “Read the text-based description of the image.” “Convert the image details into speech.” Example Usage:
[…]
  IC_STT (Speech-to-Text): “Write down what you see in the image.” “Can you write a detailed description of the picture?” “Write a summary of the scene in the image.” “Write down the elements present in the picture.” “Write down the key features visible in the image.” “In text, explain what is happening in the picture.” “Write a visual interpretation of the image.” Example Usage:
[…]
Figure 3: Prompts for Image Captioning Tasks.
VQA (Visual Question Answering) Prompts: VQA_TTT (Text-to-Text) and VQA_STS (Speech-to-Speech): Example Usage:
[…]
  VQA_STT (Speech-Text-to): “Please provide your answer in writing.” “Respond to the question with a written explanation.” “Answer this question using text.” “Type your response to the question.” “Write down your answer clearly.” “Provide a detailed answer in text format.” “Explain your response in written form.” “Answer the question by typing a full response.” “Give your answer in written words.” “Respond with a written explanation and elaborate where necessary.” “Use text to explain your answer.” “Please write your answer instead of speaking it.” “Type out your answer using complete sentences.” “Deliver your response in text form.” “Provide a clear and concise written answer.” Example Usage:
[…]
-ion using a single word or phrase."
  VQA_TTS (Text-to-Speech): ‘Respond to this question out loud.” ‘Please give your answer verbally.” ‘Provide a spoken response to the question.” ‘Answer this question using speech.” ‘Explain your response in spoken form.” ‘Say your answer instead of writing it down.” ‘Provide a verbal explanation for the question.” ‘Speak your answer clearly.” ‘Give a detailed spoken answer to this question.” ‘Respond verbally and expand on your answer.” ‘Use spoken words to explain the answer.” ‘Speak your response rather than typing it.” ‘Answer the question aloud with full sentences.” ‘Deliver your response in speech format.” Example Usage:
[…]
-ion using a single word or phrase."
Figure 4: Prompts for VQA Tasks.
A.2 Training Strategy
Our model is trained in three sequential phases, progressively increasing multimodal complexity to enhance stability, efficiency, and modality integration.
Stage 1 - Text-Speech Pre-Training
The model learns text-speech alignment through ASR and TTS on large-scale paired datasets, establishing a strong linguistic foundation before incorporating vision.
Stage 2 - Vision-Text-Speech Pre-Training
Training expands to include vision-based tasks like VQA and image captioning alongside ASR and TTS, enabling the model to unify vision, text, and speech representations.
Stage 3 - Supervised Fine-Tuning
The model is refined through supervised fine-tuning on benchmark-aligned datasets, focusing on multi-turn conversations, history retention, and multimodal reasoning.
Appendix B Implementation Details
Table 3: Implementation Details for Model, Pre-Training, and Supervised Fine-Tuning
Model and Hardware
LLM Qwen2.5-1.5B
Vision Encoder openai/clip-vit-large-patch14-336
Speech Encoder SpeechTokenizer (1s $\to$ 50 tokens),
Speech Decoder SoundStorm-SpeechTokenizer
Frameworks PyTorch, DeepSpeed
Training Configurations
Pre-Training Supervised Fine-Tuning
Batch Size 32 8
Epochs 1 2
Optimizer Adam Adam
Learning Rate $2\times 10^{-5}$ $1\times 10^{-5}$
Warmup Ratio 0.03 0.02
LR Scheduler Cosine Cosine
Training Steps 1.2M 600K
Training Epochs 1 2
Maximum Tokens 4096 4096
GPUs 8 x NVIDIA H100 4 x NVIDIA H100
DeepSpeed Config zero2
(a) Pre-training Loss
(b) SFT loss
Figure 5: Comparison of Pre-training and SFT Loss Curves
Figure 5 compares the loss curves for pre-training and supervised fine-tuning (SFT). Figure 5(a) plot shows the pre-training loss, where the raw loss exhibits high variance but progressively decreases, stabilizing after around 400K steps. The smoothed loss curve (window=90) highlights a consistent downward trend, indicating stable convergence. The 5(b) illustrates the SFT loss, comparing models with and without text-output instructions. While both configurations show a decreasing trend, the model trained with text-output instructions (red) achieves consistently lower loss than the version without instructions (green), suggesting that structured textual guidance improves fine-tuning efficiency and convergence. The variance in SFT loss remains higher compared to pre-training, reflecting the increased complexity of supervised instruction tuning.
(a) Image Caption Performance
(b) VQA Performance
(c) Overall Audio Similarity Performance
(d) Speech Image Caption Similarity
(e) Speech VQA Similarity
(f) Weighted Speech VQA Similarity
Figure 6: Performance trends of image captioning and VQA tasks over training steps.
According to Figure 6, the performance for both text and speech outputs consistently improves with additional training steps, but notable differences emerge between modalities. In text-based evaluations (Figures 6(a)6(b)), accuracy increases quickly and stabilizes after approximately 400K steps, with instruction-tuned models maintaining a moderate advantage over non-instruction models. In contrast, speech-based outputs (Figures 6(c)6(f)), evaluated via acoustic similarity, show a more pronounced improvement from instruction tuning across training steps. The performance gap between instruction-tuned and non-instruction models remains significant throughout, highlighting the critical role of instructions in enhancing speech quality and consistency. Overall, while text outputs rapidly reach a performance plateau, speech outputs benefit more substantially from extended training and explicit instruction tuning.
Appendix C Evaluation Details
Generated Modality Text Speech
Beam size 5 1
Top-P - 0.7
maximum 256 1024
Repetition Penalty 1.0 1.3
Table 4: Comparison of Text and Speech Generation Settings in evaluation.
ASR Prompt:
"Please convert this audio to text: <speech_start>{speech tokens}<speech_end>."
TTS Prompt:
"Please convert this text to speech: {transcript}."
IC_TTT and IC_STS Prompt:
[…]
IC_TTS Prompts:
[…]
IC_STT Prompts:
[…]
VQA_TTT and VQA_STS Prompt:
[…]
VQA_TTS Prompts:
[…]
using a single word or phrase."
VQA_STT Prompts:
[…]
using a single word or phrase."
Figure 7: Evaluation Prompts for different tasks.
Appendix D Examples
(a) SVLA-2B-Text-Ins
(b) SVLA-2B
Figure 8: Comparison of SVLA-2B-Text-Ins and SVLA-2B in Multimodal Image Captioning and VQA Responses (Example 2).
(a) SVLA-2B-Text-Ins
(b) SVLA-2B
Figure 9: Comparison of SVLA-2B-Text-Ins and SVLA-2B in Multimodal Image Captioning and VQA Responses (Example 3).